Qd =8 - P. Qs = 2P + 2. What is the equilibrium price and quantity?

a. P* = 2, Q* = 6

b. P* = 2, Q* = 4

c. P* = 3, Q* = 6

d. P* = 4, Q* = 8
Answer #1

Option A is correct - P* = 2 , Q* = 6

Equilibrium price and quantity can be calculated by equating the demand and supply equation.

Qd = Qs

8 - P* = 2P* + 2

6 = 3P*

P* = 2 (Equilibrium price)

Now by putting the value of equilibrium price in either of the equations of demand or supply we will get the equilibrium quantity. Substituting the value in the demand equation we get,

Qd = 8 - P*

Q* = 8 - 2

Q* = 6 (Equilibrium quantity)